OpenStudy (anonymous):

Name the property the equation illustrates. -2.1*1+ -2.1

OpenStudy (anonymous):

IS IT Identity Property of Multiplication OR Inverse Property of Multiplication

OpenStudy (skullpatrol):

Is that suppose to be $$-2.1*1= -2.1$$ not -2.1*1+ -2.1

OpenStudy (anonymous):

NO ITS RIGHT HOW I PUT IT -2.1x1=-2.1

OpenStudy (anonymous):

I gave you like all the propertied yesterday.

OpenStudy (anonymous):

and a link

OpenStudy (anonymous):

here take a look at a post you did yesterday. http://openstudy.com/users/anaivillegas#/updates/515b4223e4b0507ceba30f85

OpenStudy (anonymous):

Commutative Property of Addition a + b = b + a ex:) 2 + 3 = 3 + 2

OpenStudy (anonymous):

Commutative Property of Multiplication a • b = b • a ex:) 2 • ( 3 ) = 3 • ( 2 )

OpenStudy (anonymous):

Associative Property of Addition a + ( b + c ) = ( a + b ) + c ex:) 2 + ( 3 + 4 ) = ( 2 + 3 ) + 4

OpenStudy (anonymous):

Associative Property of Multiplication a • ( b • c ) = ( a • b ) • c ex:) 2 • ( 3 • 4 ) = ( 2 • 3 ) • 4

OpenStudy (anonymous):

Distributive Property a • ( b + c ) = a • b + a • c ex:) 2 • ( 3 + 4 ) = 2 • 3 + 2 • 4

OpenStudy (anonymous):

Additive Identity Property a + 0 = a ex:) 3 + 0 = 3

OpenStudy (anonymous):

Multiplicative Identity Property a • 1 = a ex:) 3 • 1 = 3

OpenStudy (anonymous):

Additive Inverse Property a + ( -a ) = 0 ex:) 3 + (-3) = 0

OpenStudy (anonymous):

Zero Property a • 0 = 0 ex:) 5 • 0 = 0

OpenStudy (skullpatrol):

@anaivillegas This is what you posed as a question: Name the property the equation illustrates. -2.1*1+ -2.1

OpenStudy (skullpatrol):

This is not an "equation," can you tell me why?

OpenStudy (anonymous):

First she said: -2.1*1+ -2.1 Then: -2.1x1=-2.1

OpenStudy (anonymous):

srry @skullpatrol

OpenStudy (anonymous):

yeah I see @skullpatrol haha poor you :( lol

OpenStudy (anonymous):

lol

OpenStudy (skullpatrol):

Identity Property of Multiplication: There is a unique real number 1 such that for every real number a, a*1=a and 1*a=a.

OpenStudy (skullpatrol):

Replace "a" with the real number -2.1

OpenStudy (anonymous):

or look at my example :)